• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 16:25:17 on September14,2009

    Entry number 289535

    keyword=day shift summary

    Day shift summary
    Shift worksers: J. P. Chen (SL), X. Yan (TO), K. Allada (3rd)

    8:00 smooth production data taking with 100 uA
    9:44 MCC took beam away to tune, beam was on the edge of scraping.
    (supposed to be only 10 minutes work).
    10:18 beam back, resume production. Compton was still off.
    10:34 Compton HV could not be turned on. The tuning was not good
    for the Compton. Need to retune. Stopped production run .
    11:39 MCC un-did the tuning. We are abck to what we were (before 9:44). Beam back and resum production.
    11:39-14:44 data taking
    14:44-16:00 Right Q1 fluctuation, access.
    16:00 beam tuning

    Production runs on LH2:
    12301 7:35-8:31, 100 uA, HWP IN, 100K.
    12302 8:31-9:28, 100 uA, HWP IN, 100K.
    12303 9:29-9:44, 100 uA, HWP IN, 26K, Stopped for beam tuning.
    12304 10:18-10:34, 100 uA, HWP IN, 27K. No Compton. Stopped for re-tuning for Compton.
    12305 11:39-12:36 , 100 uA, HWP IN,100K
    12306 12:36-13:33, 100 uA, HWP IN, 100K
    12307 13:33-13:54, 100uA, HWP IN, 36K, left Q1 current fluctuation
    12308, left Q1 fluctuate, junk.
    12309, 14:01-14:44, 100uA, HWP in, 75K, left Q1 still fluctuate.